a person requires 2870 kcal of energy to lead normal daily life.if heat of combustion of cane sugar is -1349 kcal then his daily consumption of sugar is 
a)728g
b)0.728 g
c)342g
d)0.342g

Grade:11

2 Answers

Vikas TU
7 years ago

C6H12O6 + 9O2→ 6CO2 + 6H2O

 
This is the heat of combustion reaction for sugar.
For 1 mol it is -1349 Kcal.
For 2870 Kcal total moles would be => 2870/1349 => 2.127 moles.
Weight of sugar => 2.127*180 = 382 gms
Dia
6 years ago
Molar mass of cane sugar (C12H22O11) is 342.2965 g/molthe heat of combustion reaction for sugar.For 1 mol it is -1349 Kcal.For 2870 Kcal total moles would be = 2870/1349 = 2.127 moles. Weight of sugar = 2.127*342. = 728 g